# TALC: Time-Aligned Captions for Multi-Scene Text-to-Video Generation

### v3

@@ 3.2 TALC Training
To this end, we propose a method for creating synthetic real-world multi-scene video-text examples, with which we further train our T2V model. Formally, we start with a video-text dataset $\mathcal{M}=\mathcal{A}\times\mathcal{B}$ consisting of pairs of $(A_{i},B_{i})$ where $A_{i}$ is a raw video and $B_{i}$ is the corresponding video description. Then, similar to [9], we apply PySceneDetect [^2] https://github.com/Breakthrough/PySceneDetect to segment $A_{i}$ into continuous video scenes $A_{i}=\{A_{i,1},A_{i,2},\ldots,A_{i,m}\}$ where $m$ is the number of detected scenes. Next, we sample the middle video frame $F_{i,j}$ in each scene $A_{i,j}$ as its semantic representative. Finally, we input all the video frames $F_{i}=\{F_{i,1},\ldots,F_{i,m}\}$ for $A_{i}$ and the entire video caption $B_{i}$ to Gemini-Pro-Vision-1.0 [38]. Gemini is prompted to generate captions for each of the frames $F_{i,j}$ such they form a coherent narrative guided by the common caption $B_{i}$. In Figure 3, we provide an instance for the multi-scene video-text data generation. We provide the prompt provided to Gemini and the source datasets in Appendix §D and §E, respectively.
4 Experimental Setting
In this section, we describe the setup of our experiments for generated multi-scene videos, including eval-set creation, metrics, automatic and human evaluation, and the tested models.
4.1 Eval-Set Task Prompts
To create an eval-set for testing models that generate multi-scene videos, we curate a set of task prompts for diverse scenarios (The full task prompts are in Appendix §F.). We generate $100$ prompts for each scenario.
Single character in multiple visual contexts (S1): We instruct GPT-4 [39] to create a coherent script consisting of four scenes. Each scene features a specific animal character performing diverse activities in every scene. This task assesses the consistent appearance of the entity and its background while adhering to the different actions (or events) described in the multi-scene text script. An example script is “Scene 1: A red panda is climbing a tree. Scene 2: The red panda eats the leaves on the tree. Scene 3: … Scene 4: …”.
Different characters in a specific visual context (S2): We instruct GPT-4 to create a coherent script consisting of four scenes. Each scene features the same activity but with different animal characters [18]. This task assesses the capability of the T2V model to generate consistent appearance of the background while adhering to the animal character changes. An example script is “Scene 1: A cat leaps onto a countertop. Scene 2: A dog leaps onto the same countertop. Scene 3: A rabbit … Scene 4: A raccoon …”.
Multi-scene captions from real videos (S3): We use our multi-scene video/text data generation pipeline (§3.2) to create similar prompts for real videos for the test splits of the video-text datasets. This task assesses the ability of a model to generate multi-scene videos with consistent appearances of entities and background for open-ended prompts, which are derived from real videos. An example script is “Scene 1: A vlogger introduces her skincare routine. Scene 2: She applies a serum to her face. Scene 3: …”.
4.2 Evaluation Procedure
Following [27], we assess the capability of the tested T2V models to: (a) follow the text instructions in the script (Text Adherence); (b) keep consistent appearance of entities and background across scenes (Visual Consistency), unless the script specifically requests a change in any of them.
@@ 4.2 Evaluation Procedure
We compute the visual consistency and text adherence with automatic and human evaluators. We also report an overall score, which is the average of the visual consistency and text adherence scores. In addition, we also assess the visual quality of the generated videos using human evaluation to understand whether the video contains any flimsy frames, shaky images, or undesirable artifacts (Appendix M):
Automatic Evaluation: We utilize GPT-4-Vision (GPT4V) [40] to reason over multiple images. First, we sample four video frames, uniformly, from each scene in the generated video (e.g., 8 videos frames for a two-scene video). Then, we prompt GPT4V with the temporal sequence of the sampled frames and the multi-scene text description and instruct GPT4V to decide on the quality of the generated video across several attributes, including entity consistency, background consistency, and text adherence. For each attribute, GPT4V assigns one of three possible response $\{\textit{yes}=1,\textit{partial}=0.5,\textit{no}=0\}$. The automatic evaluation prompt is given in Appendix §G.
[…]
@@
Since merging captions (§4.3.1) and TALC3.1) methods input the entire multi-scene text description at once, the quality of the video generated by these methods is influenced by the number of scenes described in the text script. Hence, we calculate the performance of the tested models by averaging the scores assigned to videos generated for two, three, and four scenes, over 900900900900333100 (examples per scenario) * 3 (scenarios) * 4 (2-scene/3-scene/4-scene) = 900. multi-scene prompts. We report visual consistency by averaging the performance across the entity and background consistency metrics. Here, the entity consistency scores are calculated for the task prompts S1 and S3 (since S2 aims to change the characters across scenes), while the background consistency and text adherence are computed for all the task prompts.
4.3 Tested Models
We experiment with ModelScope [10] and Lumiere [11] T2V models for multi-scene video generation with TALC. We mainly experiment with ModelScope due to its ease of reproduction and adoption in prior work [27]. We also experiment we Lumiere to showcase TALC’s generalization.
By design, TALC can be applied to a base T2V model during inference, without any multi-scene finetuning. We compare the multi-scene performance of the two base models without finetuning under three settings: TALC, merging captions, and merging videos (see baseline details in Section 4.3.1).
In addition, we finetune ModelScope with the TALC framework using the multi-scene video-text data (§3.2). As a baseline, we also finetune ModelScope without the TALC framework by naively merging the scene-specific captions in the raw text space. In both cases, we finetune the T2V model with $8$ frames per scene and the maximum number of scenes in an instance is set to $4$. We provide more finetuning details in Appendix §L.
During inference, we generate $16$ frames per scene from ModelScope and $80$ frames per scene from Lumiere. More inference details are found in Appendix §J.
4.3.1 Baselines
For Merging Captions, we create a single caption by merging all the multi-scene descriptions. Specifically, the multi-scene descriptions $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ can be written as a single prompt ‘$\mathcal{P}=y_{1}.\text{Then, }y_{2}.\ldots\text{Then, }y_{n}.$’. For instance, the two-scene description $Y=$‘A red panda climbs on a bamboo forest.’, ‘The red panda sleeps peacefully in the treetop.’} would change to $\mathcal{P}=$A red panda climbs on a bamboo forest. Then, the red panda sleeps peacefully in the treetop.” Subsequently, we generate a video from the T2V model $f_{\theta}$ by conditioning it on $\mathcal{P}$.
For Merging Videos, we generate a video $x_{i}$ for each scene description $y_{i}$ individually and stitch the output videos together afterwards to synthesize the entire video $X={x_{1},x_{2},\ldots,x_{n}}$.
@@ 5 Experiments
(a) Performance on ModelScope T2V model.
(b) Performance on Lumiere T2V model.
Figure 4: Automatic evaluation results for (a) ModelScope and (b) Lumiere. In (a), we observe that TALC-finetuned ModelScope model achieves the highest overall score, that is the average of the visual consistency and text adherence scores. In (b), we find that TALC framework with the Lumiere base model outperforms merging captions and merging videos on the overall scores. We report the average performance across the diverse multi-scene prompts and the number of generated scenes.
5.1 Automatic Evaluation Results
@@
We compare the performance of the baselines (e.g., merging captions and merging videos) with the TALC framework for ModelScope and Lumiere using the automatic evaluation in Figure 4.
@@
In Figure 4(a), we find that the overall score, average of visual consistency and text adherence, of the multi-scene videos generated using the base ModelScope with TALC ($69$ points), outperforms the overall score achieved by the videos generated using merging captions ($62$ points) and merging videos ($68$ points) with the base ModelScope. Specifically, we observe that the visual consistency of the generated video is high for merging captions ($91$ points) and TALC ($89.9$ points) while it is low for merging videos ($65$ points). This indicates that merging videos independently for the individual scene descriptions does not preserve the background and entity appearances across the different frames.
Since merging captions mentions multiple events in a single prompt, the T2V model struggles to understand the temporal boundaries between these events. As indicated by Eq. 2, the visual features for all video frames aggregate information from the entire multi-scene description simultaneously, without aligning individual scene descriptions with their corresponding appearances in the generated video. In Figure 4(a), this results in poor text adherence with merging captions ($32$ points).
In addition, we observe that the text adherence using TALC outperforms merging captions by $14.8$ points, while the text adherence is the highest with a score of $70$ points using merging videos. This can be attributed to the design of the merging videos baseline where individual video scenes adhere to the scene-specific descriptions well. Hence, merging videos independently approach can be viewed as an upper bound on the text adherence metric.
In Figure 4(b), we observe similar trends for the Lumiere T2V generative model. Specifically, we find that the overall score for TALC outperforms merging captions and merging videos by $4$ points and $2$ points, respectively. In addition, we observe that merging captions and TALC achieve a high visual consistency score while merging videos independently has poor visual consistency. Further, we find that TALC outperforms merging captions by $5$ points on text adherence, while merging videos achieves the highest text adherence $65$ points. This highlights that the TALC framework enhances the adherence to the scene descriptions.
@@
Earlier, we evaluated the usefulness of the TALC framework with the base model. However, the base models are trained with the single-scene video-text data that might limit their capability for multi-scene video generation. To alleviate this issue, we finetune ModelScope T2V model on the multi-scene video-text data (§3.2). Specifically, we finetune the model using the merging captions method and TALC framework, independently.
In Figure 4(a), we find that finetuning with TALC achieves the highest overall score of $75.6$ points in comparison to all the baselines. Specifically, we observe that the visual consistency does not change much with finetuning using the TALC method ($89.9$ points vs $89$ points). Interestingly, we observe that finetuning with merging captions reduces the visual consistency by a large margin of $14$ points. This can be attributed to the lack of knowledge about the natural alignment between video scenes and individual scene descriptions, which gets lost during the merging of captions.
Additionally, we find that the text adherence of the TALC-finetuned model is $15.1$ points more than the text adherence of the TALC-base model. Similarly, we find that the text adherence of the merging captions-finetuned model is $5.1$ points more than the text adherence of the merging captions-base model. This highlights that finetuning a T2V model with multi-scene video-text data helps the most with enhancing its text adherence capability.
@@
5.2 Human Evaluation Results
Figure 5: Human evaluation results for ModelScope model. We observe that the base model using the TALC framework outperforms the merging captions and merging videos baselines on the overall score. In addition, TALC-finetuned model enhances the text adherence and achieves the highest overall score. We report the average performance across the diverse multi-scene prompts and the number of generated scenes.
TALC achieves the best performance in human evaluation.
We compare the performance of the baselines and TALC framework for ModelScope using human evaluation in Figure 5. We find that TALC-finetuned model outperforms the merging captions and merging video methods with the base model by $12$ points and $15.5$ points, respectively, on the overall score. In addition, we find that using TALC framework in the base model outperforms the merging captions and merging video methods with the base model by $7.6$ points and $11.1$ points, respectively, on the overall score. Further, we observe that the merging captions with the base model achieves the highest visual consistency score of $96.5$ points while it is the lowest for merging videos generated from the base model. In addition, we find that the text adherence of the TALC-finetuned and TALC-base model is better than merging captions-finetuned and merging captions-base model, respectively. Our results highlight at the benefit of including the inductive bias of temporal alignment between the video scenes and their scene descriptions for multi-scene video generation. We perform human evaluation on the visual quality of the generated multi-scene videos in §M. While our work is centered around multi-scene evaluation, we also perform single-scene evaluation in Appendix §N.
@@ 5.3 Qualitative Analysis
We provide qualitative examples of generating multi-scene videos using ModelScope (fine-tuned with TALC) and Lumiere (base model with TALC) for diverse scenarios in Figure 12. Our analysis reveals that both ModelScope and Lumiere are capable of producing multi-scene videos that exhibit high text adherence and visual consistency. Considering the case of the same animal engaging in multiple actions (referred to as ‘one character multiple contexts’). The videos generated by ModelScope successfully maintained the same animal while varying the background and action between the scenes. Conversely, the videos generated by Lumiere displayed the same animal performing different actions with minimal background alterations. We believe that this distinction is attributed to ModelScope’s fine-tuning with TALC. Considering different animals within a particular visual setting (referred to as ‘multiple-characters same context’), both ModelScope and Lumiere demonstrated impressive abilities in preserving the consistency of the background across the videos and adhering closely to the provided text. During our analysis, we noticed that the multi-scene captions derived from real videos (referred to as ‘open-ended captions’) exhibited a substantial number of changes between the various scenes. In this scenario, Lumiere, when employed without fine-tuning, displayed challenges in adhering to the text, while ModelScope achieved a higher degree of text adherence but was also prone to visual artifacts.
6 Related Work
Text-to-Video Generative Modeling.
The field of text-to-video (T2V) synthesis has significantly evolved from its inception with models like VGAN [2] and MoCoGAN [41], leveraging the foundational technologies of GANs [42] and VAEs [43] to produce concise, single-scene videos. The advent of diffusion models, exemplified by Imagen Video [44], marked a notable advancement in T2V synthesis. Despite these strides, the challenge of creating multi-scene videos that reflect the complexity of the physical world [1, 2, 3] remains. Our work, TALC, extends the capabilities of T2V models to multi-scene storytelling, filling a crucial gap in the synthesis landscape.
Multi-Scene Video Generation.
The pursuit of multi-scene T2V synthesis has been furthered by recent innovations like Phenaki [18] and Stable Video Diffusion [9], which have explored new frontiers in video generation from textual prompts and the scaling of latent diffusion models, respectively. Despite these advancements, the task of generating videos that convey coherent narratives across multiple scenes remains formidable, highlighted by recent works such as VideoPoet [19]. TALC tackles this task and offers a framework produces videos spanning multiple scenes. We also introduce nuanced evaluation approach. This approach integrates both automated assessments and human evaluations to rigorously gauge the quality and narrative coherence of the generated content, evaluating text adherence, object consistency and background consistency, contributing to the ongoing refinement of T2V synthesis. We provide a detailed related work in §B.

### v4

@@ 3.2 Multi-Scene Video-Text Data Generation
While our approach generates better multi-scene videos, the text adherence capabilities of the pretrained T2V generative model are limited. This is due to the lack of multi-scene video-text data during its pretraining. To this end, we create a real-world multi-scene video-text dataset to allow further training of the pretrained T2V models. Specifically, we leverage the capability of the multimodal foundation model, Gemini-Pro-Vision [43], to generate high-quality synthetic data for enhanced video-text training [5]. Formally, we start with a video-text dataset $\mathcal{M}=\mathcal{A}\times\mathcal{B}$ consisting of pairs of $(A_{i},B_{i})$ where $A_{i}$ is a raw video and $B_{i}$ is the corresponding video description from the dataset. Subsequently, we utilize PySceneDetect library  [^1] https://github.com/Breakthrough/PySceneDetect to generate continuous video scenes from $A_{i}=\{A_{i,1},A_{i,2},\ldots,A_{i,m}\}$ where $m$ is the number of scene cuts in the video. Further, we sample the middle video frame $F_{i,j}$ as a representative of the semantic content in the video scene $A_{i,j}$. Finally, we input all the video frames $F_{i}=\{F_{i,1},\ldots,F_{i,m}\}$ for a single video $A_{i}$ and the entire video caption $B_{i}$ to Gemini-Pro-Vision. Specifically, the model is prompted to generate high-quality captions for each of the frames $F_{i,j}$ such they form a coherent narrative guided by the common caption $B_{i}$. We provide the prompt provided to the multimodal model in Appendix §E. In Figure 4 we provide an instance for the multi-scene video-text data generation. We highlight that higher-quality multi-scene datasets would enhance the performance of the models with TALC framework. We provide the more details in Appendix C.  [^2] We note that the parts of this pipeline are also utilized in a concurrent work, ShareGPTVideo [15]. However, their approach is focused on long-form video generation pretraining, while we are focus on multi-scene video generation for finetuning.
4 Evaluation
4.1 Metrics
The ability to assess the quality of the generated multi-scene videos is a challenging task itself. As humans, we can judge the multi-scene videos across diverse perceptual dimensions [25] that the existing automatic methods often fails to capture [11]. Following [30], we focus on the visual consistency of the generated video, and text adherence capabilities of the T2V models.
Visual Consistency. This metric aims to assess the (entity or background) consistency between the frames of the multi-scene videos. Here, the entity consistency aims to test whether the entities in the multi-scene video are consistent across the video frames. For instance, the appearance of an animal should not change without a change described in the text description. In addition, the background consistency aims to test whether the background of the multi-scene video remains consistent across the video frames. For instance, the room should not change without a change in text description.
Text Adherence. This metric aims to test whether the generated video adheres to the multi-scene text description. For instance, the events and actions described in the text script should be presented in the video accurately, and in the correct temporal order.
@@ 4.1 Metrics
In our experiments, we compute the visual consistency and text adherence with the human and automatic evaluators. Further, we compute the overall score, which is the average of the visual consistency and text adherence scores. In addition, we also assess the visual quality of the generated videos using human evaluation to understand whether the video contains any flimsy frames, shaky images, or undesirable artifacts (Appendix D).
4.2 Task Prompts
Here, we curate task prompts for diverse scenarios to holistically assess the quality of the videos.
Single character in multiple visual contexts (S1). In this scenario, we instruct an LLM, GPT-4 [2], to create a coherent script consisting of four scenes. Each scene features a specific animal character performing diverse activities in every scene. This task assesses the capability of the T2V model to generate consistent appearance of the entity and its background while adhering to the different actions (or events) described in the multi-scene text script.
Different characters in a specific visual context (S2). In this scenario, we instruct a language model, GPT-4, to create a coherent script consisting of four scenes. Each scene features different animal characters engaging in the same activity in every scene [45]. This task assesses the capability of the T2V model to generate consistent appearance of the background while adhering to the appearance of the different characters in the multi-scene text script.
Multi-scene captions from real videos (S3). Here, we aim to assess the ability of the model to generate multi-scene videos for open-ended prompts that are derived from real-world videos. This task also assesses the ability of the T2V model to generate consistent appearances of the entity and its background while adhering to multi-scene descriptions. Specifically, we use our multi-scene video-text data generation pipeline (§3.2) to create such prompts for the real videos from the test splits of the video-text datasets. In total, we generate $100$ prompts in this scenario. We present a sample of the various task prompts in the Appendix §F. We present the details about the human and automatic evaluators in §B.
4.3 Evaluation Setup
@@
We compute the performance of the baselines and TALC by averaging the scores assigned to videos generated for two, three, and four scenes. Additionally, we report on visual consistency by averaging the performance across the entity and background consistency metrics. Here, the entity consistency scores are calculated for the task prompts S1 and S3 (since S2 aims to change the characters across scenes), and the background consistency and text adherence scores are computed for all the task prompts. We also evaluate the impact of TALC-based finetuning for single-scenes in Appendix §M.
Table 1: Human evaluation results. We present the human evaluation results for the overall score for several baselines (e.g., FreeNoise, merge videos, and merge captions) and TALC framework for the ModelScope generative model. Specifically, we find that the TALC with the base model outperforms the FreeNoise by achieving the relative gain of $21\%$ on the overall score. In addition, finetuning ModelScope with TALC framework enables better multi-scene video generation by achieving the highest overall score. Overall, TALC strikes a good balance between visual consistency and text adherence.
Overall score Visual consistency Text adherence
FreeNoise [38] 59.7 77.0 42.5
Merge videos 61.3 ($+2.6\%$) 55.0 67.5
Merge captions 64.8 ($+8.5\%$) 96.5 33.0
Finetuning w/ Merge captions 61.1 ($+2.3\%$) 80.0 42.3
TALC 72.4 ($+21\%$) 92.3 52.5
Finetuning w/ TALC 76.8 ($+29\%$) 86.4 67.2
@@ 5 Experiments
5.1 Text-to-Video Generative Models
We perform most of our experiments on ModelScope [49] due to its easy-of-access and adoption in prior works [30]. In addition, we also include Lumiere-T2V, a model that leverages space-time U-Net denoising networks to generate high-quality videos. In this work, we include early experiments with Lumiere to showcase the flexibility of the TALC approach on diverse models. We perform human evaluation for ModelScope, and automatic evaluation for both ModelScope and Lumiere.
Base model with TALC. As described in §3.1, our approach modifies the traditional text-conditioning mechanism to be aware of the alignment between text descriptions and individual video scenes. By design, the TALC framework can be applied to the base T2V model during inference, without any multi-scene finetuning. Here, we generate $16$ frames per scene from ModelScope and $80$ frames per scene from Lumiere. We provide more details in Appendix §J.
Finetuning with TALC. Since the base model is pretrained with single-scene data, we aim to show the usefulness of TALC framework when we have access to the multi-scene video-text data. To this end, we finetune ModelScope on the multi-scene video-text data (§3.2) with TALC framework. As a pertinent baseline, we also finetune the ModelScope without TALC framework by naively merging the scene-specific captions in the raw text space. In this setting, we finetune the T2V model with $8$ frames per scene and the maximum number of scenes in an instance is set to $4$. We provide further details on the finetuning setup in Appendix §L. The inference settings are identical to the prior method of generating videos from the base model without finetuning.
In this section, we present the results for the baselines and TALC framework averaged over a diverse task prompts and multiple scenes using automatic evaluation (§5.4) and human evaluation (§5.3). Finally, we provide qualitative examples for the multi-scene generated videos to showcase the usefulness of our approach (§5.5).
5.2 Baselines
Merge Captions
In this setup, we create a single caption by merging all the multi-scene descriptions. While this approach mentions the temporal sequence of the events in a single prompt, the T2V model does not understand the temporal boundaries between the two events. Specifically, the visual features for all the frames will aggregate information from the entire multi-scene description, at once, without any knowledge about the alignment between the scene description and its expected appearance.
Merge Videos
In this setup, we generate videos for each scene description individually and merge them in the raw input space. In this process, the parts of the multi-scene video closely adhere to the scene descriptions, leading to high text fidelity. However, since the generated videos do not have access to all the multi-scene descriptions (e.g., the video for Scene 2 is not informed about Scene 1), the visual consistency across the entire video is quite poor.
FreeNoise
We also consider a more sophisticated approach, FreeNoise [38]. Specifically, it reschedules a sequence of noises for long-range correlation and perform temporal attention over them by window-based fusion. In addition, it includes motion injection method to support the generation of videos conditioned on multiple text prompts. Due to its complexity, this method involves setting various hyperparameters, including the denoising step at which motion injection should be activated, the cross-attention layers where prompts are injected, and the frames between which the prompt representations should be interpolated. In this work, we evaluate the publicly available implementation with VideoCrafter [13] on our unseen prompts. [^3] https://github.com/AILab-CVC/FreeNoise While Gen-L-Video [48] has been shown to perform worse than FreeNoise, we conduct a qualitative analysis to emphasize the robustness of our method against it. (§5.5). [^4] Other approaches such as VideoDirectorGPT [30] and Phenaki [45] are not publicly available.
5.3 Human Evaluation
TALC achieves the best performance in human evaluation.
We compare the performance of the baselines and TALC framework using human evaluation in Table 1. We find that TALC applied to ModelScope (MS) outperforms FreeNoise by achieving a relative gain of $21\%$ on the overall score. In addition, finetuning MS with TALC framework increases the video generation capability, leading to relative gain of $27\%$ over FreeNoise in the overall score. The poor performance of FreeNoise can be attributed to the inability to introduce new content and limited motions as the multiple scenes ($>2$) are requested. We provide some qualitative examples for FreeNoise to highlight its limitations in Appendix N. In addition, we find that using TALC framework in the base model outperforms the merging captions and merging video methods with the base model by $7.6$ points and $11.1$ points, respectively, on the overall score. Further, we note that TALC-finetuned model outperforms the merging captions and merging video methods with the base model by $12$ points and $15.5$ points, respectively, on the overall score.
Additionally, we observe that the merging captions with the base model achieves the highest visual consistency score of $96.5$ points while it is the lowest for merging videos generated from the base model. Our results indicate that merging videos independently for the individual scene descriptions does not preserve the background and entity appearances across the different frames. Further, we note that the text adherence of the TALC-finetuned and TALC-base model is better than merging captions-finetuned and merging captions-base model, respectively. The high text adherence for merging videos can be attributed to its design where individual video scenes adhere to the scene-specific descriptions well. Overall, our empirical findings highlight that our simple framework can enable robust multi-scene video generation with finetuning. We present the results for the visual quality metric in Appendix D.
Table 2: Automatic evaluation results for ModelScope. We compare the performance of the baselines and TALC framework using the automatic evaluation for ModelScope generative model. Similar to the human evaluation, we observe that our simple approach achieves the best overall score with the base model (TALC) and finetuned model (Finetuning w/ TALC). We abbreviate visual consistency as VC, and text adherence as TA.
Overall score VC TA
Merge captions 61.7 91.0 32.4
Merge videos 67.5 ($+9.4\%$) 65.0 70.0
Finetuning w/ Merge captions 57.3 ($-7.1\%$) 77.0 37.5
TALC 68.6 (+11.2%) 89.9 47.2
Finetuning w/ TALC 75.6 (+22.5%) 89.0 62.3
Table 3: Automatic evaluation results for Lumiere. We present the results for the comparison between the baselines and TALC for the Lumiere video generative model. Specifically, our results indicate that TALC achieves the highest overall score, by achieving $7.1\%$ relative gains over merge captions on the overall score. This indicates that TALC is a flexible strategy that can be applied to diverse video generative models. We abbreviate visual consistency as VC, and text adherence as TA.
Overall score VC TA
Merge captions 64.4 94.7 34
Merge videos 66.5 (+3.2%) 68.0 65.0
TALC 69.0 (+7.1%) 97.8 40.0
5.4 Automatic Evaluation
@@
We compare the performance of the baselines with the TALC framework for ModelScope and Lumiere using the automatic evaluation in Table 2 and 3, respectively.
@@
In Table 2, we find that the overall score, average of visual consistency and text adherence, of the multi-scene videos generated using the base ModelScope with TALC ($68.6$ points), outperforms the overall score achieved by the videos generated using merging captions ($61.7$ points) and merging videos ($67.5$ points) with the base ModelScope. In addition, we observe that the text adherence using TALC outperforms merging captions by $14.8$ points, while the text adherence is the highest with a score of $70$ points using merging videos. In Table 3, we observe similar trends for the Lumiere T2V generative model. Specifically, we find that the overall score for TALC outperforms merging captions and merging videos by $4$ points and $2$ points, respectively. In addition, we observe that merging captions and TALC achieve a high visual consistency score while merging videos independently has poor visual consistency. Further, we find that TALC outperforms merging captions by $5$ points on text adherence, while merging videos achieves the highest text adherence $65$ points. This highlights that the model more easily generates multi-scene videos that adhere to individual text scripts, whereas adherence to the text diminishes when the model is given descriptions of multiple scenes all at once.
@@
In Table 2, we find that finetuning with TALC achieves the highest overall score of $75.6$ points in comparison to all the baselines. Specifically, we observe that the visual consistency does not change much with finetuning using the TALC method ($89.9$ points vs $89$ points). Interestingly, we observe that finetuning with merging captions reduces the visual consistency by a large margin of $14$ points. This can be attributed to the lack of knowledge about the natural alignment between video scenes and individual scene descriptions, which gets lost during the merging of captions. Additionally, we find that the text adherence of the TALC-finetuned model is $15.1$ points more than the text adherence of the TALC-base model. This highlights that finetuning a T2V model with multi-scene data helps the most with its text adherence capability.
@@ 5.5 Qualitative Analysis
We provide qualitative examples of generating multi-scene videos using TALC, FreeNoise and Gen-L-Video in Appendix Figure 10 and Figure 11. Our analysis demonstrates that all methods are capable of generating multi-scene videos that exhibit a high degree of text adherence. However, the primary distinction lies in the quality of the videos. FreeNoise is capable of producing videos characterized by superior quality and visual coherence. Nonetheless, the motion within these videos is notably limited, resulting in a relatively static scene. Gen-L-Video can generate videos that incorporate motion. However, visual consistency is not maintained throughout the video. For instance, in the provided example depicting a man engaged in both surfing and skiing, the man’s appearance undergoes noticeable changes across the video. Overall, we observe that TALC is capable of generating realistic multi-scene videos that not only exhibit high textual adherence but also maintain visual consistency. Furthermore, the transitions between scenes are smooth and natural.